W.O for 2362 C being too cold. Check area, 65 degrees. Also noticed hallways were cold. Check mer 13 t.u,. commanded off. System Working as programmed.
The cooling tower for chiller 4 & the yard towers for chillers 1&2 have had the float valves removed and are under Metasys control (modulating ball valve near the make-up water meter). The tower level reading is the distance from the water surface to the ultrasonic sensors face. I have decided to use the default settings so, ten inches is the min & forty inches is the max range. Just by luck, both setpoints ended up at 15″. The new points can be found under chiller #4 & yard towers, setpoints will be maintained weather it’s running or not. There will be a enable\disable point for the operator in the future.
Found out through a bit of testing with a VANEOMETER; the door separating the control room section from the procedure area of Surgery’s O.R. Hybrid Room #1 needs to be closed to maintain proper room pressurization. Thanks Cori for your efforts overnight ~ especially today… We’re under inspection for the whole OSC by JCAHO. It would not have been favorable for them to find this room in alarm, even if this is a FAKEO JCAHO!
Copied from an email notfication sent from Paul Dobry earlier:
Hello: I wanted to make everyone aware that John Darr Mechanical (JDM) will be working on a punch list item in Mechanical Equipment Room #32 serving new Outpatient Surgery Center (OSC). At or near 4 AM on Tuesday (8/26/14) JDM will secure power to the pumps serving the Terminal Unit (TU) system to perform some required re-piping. During this time no hot water will be available for reheat throughout OSC. SE on-duty should expect Metasys alarms for pump command not matching status. Pumps will be re-started by 6 AM. Bryan Diuble will be the point person for JDM and can be reached on his mobile phone if necessary at 734-216-5444.
Please let me know if you have any questions or concerns. Feel free to distribute to your teams as necessary. THANKS!
Kyle Morrison from JCI will be here tomorrow morning at or around 6 AM to replace a faulty pressure sensor serving our HYBRID OPERATING ROOM. (#2850) As is now, the room has been overridden to maintain positive pressure conditions. When finished, this will allow both s.a. and r.a. to modulate as needed.
Surgery staff have been notified, and since cases are delayed on Tuesdays anyway until 8:30, there should be no issues.